solve the system by elimination
5x+5y= -13
7x-3y= 11
use integers or fractions

Answer by checkley77(12569)

5x+5y= -13 MULTIPLY BY 3
7x-3y= 11 MULTIPLY BY 5 & ADD.
15X+15Y=-39
35X-15Y=55
--------------------
50X=16
X=16/50
X=8/25 ANS.
5*8/25+5Y=-13
40/25+5Y=-13
5Y=-13-8/5
5Y=(-13*5-8)/5
5Y=(-65-8)/5
5Y=-73/5
Y=(-73/5)/5
Y=-73/5*1/5
Y=-73/25 ANS.
PROOF:
7(8/25)-3(-73/25)=11
56/25+219/25=11
275/25=11
11=11